# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ 3.3 Bayesian model comparison
$\text{DIC}_{1}=-4E_{\boldsymbol{\phi}}[\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})]+2~{}\text{log}~{}p(\textbf{y}\mid\overline{\boldsymbol{\phi}}),$ (18)
where $\overline{\boldsymbol{\phi}}$ is the posterior mean of ${\boldsymbol{\phi}}$, and $p(\textbf{y}\mid\boldsymbol{\phi})$ denotes the density of y given the model parameters $\boldsymbol{\phi}$. Both the expectation and the posterior mean $\overline{\boldsymbol{\phi}}$ are computed using samples drawn from the approximate posterior distribution estimated by the proposed VB method. For all SEMs, the logarithm of the density $p(\textbf{y}\mid\boldsymbol{\phi})$ is provided in Section S5 of the online supplement.
@@
$\text{DIC}_{2}=-4E_{\boldsymbol{\phi}}[\text{log}~{}p(\textbf{y}\mid\boldsymbol{\phi})]+2~{}\text{log}~{}p(\textbf{y}\mid\widehat{\boldsymbol{\phi}}),$ (19)
where $\widehat{\boldsymbol{\phi}}$ is the value of ${\boldsymbol{\phi}}$ that maximises the function $p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, estimated using samples drawn from the approximate posterior obtained by the proposed VB method. The expectation is also computed using these variational posterior samples.
@@
$\text{DIC}_{5}=-4E_{\boldsymbol{\phi},\boldsymbol{\psi},\textbf{y}_{u}}[\text{log}~{}p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})]+2~{}\text{log}~{}p(\textbf{y}_{o},\widehat{\textbf{y}}_{u},\textbf{m}\mid\widehat{\boldsymbol{\phi}},\widehat{\boldsymbol{\psi}}),$ (20)
where $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})$ is the joint density of y and m, conditional on model parameters $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$. The values $\widehat{\boldsymbol{\phi}}$, $\widehat{\boldsymbol{\psi}}$, and $\widehat{\textbf{y}}_{u}$ are the posterior sample values of $\boldsymbol{\phi}$, $\boldsymbol{\psi}$ and $\textbf{y}_{u}$ that maximise the density $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})p(\boldsymbol{\phi},\boldsymbol{\psi})$, where the samples are drawn from the approximate posterior estimated by the proposed HVB method. The expectation is also computed using these variational posterior samples. Note that $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$; see the selection model factorisation in Equation (11).
4 Simulation study
The simulation study is divided into two subsections. Section 4.1 compares the posterior distributions obtained using variational Bayes (VB) methods with those of Hamiltonian Monte Carlo (HMC, Neal, (1996)). The posterior densities estimated using the HMC method are considered as ground truth to assess the accuracy of the VB methods. In Section 4.1, we utilise a small simulated dataset ($n=625$), as implementing HMC is computationally intensive for SEMs, particularly when handling large datasets with missing data (Wijayawardhana et al.,, 2025). Section 4.2 uses two moderately large simulated datasets, each considered with and without missing values, and each exhibiting distinct characteristics, to evaluate the robustness of the proposed models and VB methods.
We now describe the prior distributions used in this study for both the VB and HMC methods, applied in the simulation and real-world examples presented in Sections 4.1, 4.2, and 5. To map the parameters $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$ onto the real line, we apply the following transformations: $\omega^{\prime}=\text{log}(\sigma^{2}_{\textbf{e}})$, $\rho^{\prime}=\text{log}(1+\rho)-\text{log}(1-\rho)$, $\nu^{\prime}=\text{log}(\nu-3)$, and $\gamma^{\prime}=\text{log}(\gamma)-\text{log}(2-\gamma)$. Table S3 gives the prior distributions for the transformed parameters.
@@ 4 Simulation study
Table 3: Prior distributions of model parameters
4.1 Assessing the accuracy of variational Bayes approximations
This section compares the accuracy of the proposed VB approximation methods with the computationally intensive, exact HMC methods. The comparison focuses solely on the YJ-SEM-Gau model, both with and without missing data. To facilitate this comparison, a dataset with 625 observations is simulated from the YJ-SEM-Gau model with five covariates generated from standard normal distributions. The weight matrix, W is constructed based on a $25\times 25$ regular grid, where neighbouring units are defined using the rook neighbourhood criteria, as described in Lloyd, (2010).
@@ 4.1 Assessing the accuracy of variational Bayes approximations
The model parameters are specified as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from discrete uniform values between $-3$ and $3$ (excluding 00). The error variance parameter is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autocorrelation parameter is set to $\rho=0.8$. The Yeo Johnson (YJ) parameter is specified as $\gamma=1.25$, inducing moderate left skewness into the simulated response variable. See Figure S1 in Section S7.1 of the online supplement for the kernel density plot of the response variables.
@@
The VB algorithm, outlined in Algorithm S1 in Section S2 of the online supplement, and the HMC algorithm are used to fit the YJ-SEM-Gau model to the simulated dataset. The initial values for the VB algorithm are set as follows: The variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau model via maximum likelihood estimation (MLE) method. The initial value for $\gamma$ is set to 1. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. We use $p=4$ factors. The results do not improve when we increase the number of factors. We used similar starting values for the HMC algorithms as well.
Both the VB and HMC algorithms are run for 10,000 iterations, at which point convergence is achieved. For the HMC algorithm, the first 5,000 iterations are treated as burn-in and discarded. Convergence for the VB algorithm is assessed visually by inspecting the plots of variational means over iterations. For the HMC algorithms, convergence is assessed by inspecting trace plots of the model parameters (see Section S9.1 of the online supplement).
@@
The HMC algorithm directly generates posterior samples of the model parameters. In contrast, the VB algorithm first estimates the variational parameters, $\boldsymbol{\lambda}$, by running the variational optimisation. Then, 10,000 posterior draws of the model parameters are generated from the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$.
Figure 1 compares the posterior densities of selected model parameters estimated using the VB method and the HMC algorithm implemented via RSTAN (Stan Development Team,, 2023) to fit the simulated dataset. For the fixed effect parameter (${\beta}_{1}$), the spatial autocorrelation parameter ($\rho$), the variance parameter ($\sigma^{2}_{\textbf{e}}$), and the YJ parameter $\gamma$, the posterior densities estimated by both methods are nearly identical and include the true values.
@@
Figure 1: Comparison of posterior densities of selected parameters in the YJ-SEM-Gau model, obtained using the HMC and VB methods on the simulated dataset without missing values. The vertical line represents the true parameter value.
@@
Next, we introduce missing values into the response variable y under the MNAR mechanism, using the logistic regression model described in Section 2.3. To do this, we generate an additional covariate $\textbf{x}^{*}$ from a log-normal distribution with a log-scale mean of 0 and a log-scale standard deviation of 1. This covariate is included in the missing data model, which is specified as a logistic regression model with $\textbf{x}^{*}$ and the response variable y as covariates (see Equation (10)). The regression coefficients are set as follows: $\psi_{0}=-1.0$ (intercept), $\psi_{\textbf{x}^{*}}=0.5$ (the coefficient of $\textbf{x}^{*}$), and $\psi_{\textbf{y}}=-0.1$ (the coefficient of y). This configuration yields approximately $50\%$ missing values in y, corresponding to $n_{u}=320$.
The HVB-NoB algorithm (without block-wise update) described in Algorithm 1 and the HMC algorithm (implemented via RSTAN) are used to estimate the YJ-SEM-Gau model. The tuning parameter, $N_{1}$ of the HVB-NoB algorithm, is set to 10. The initial values for the algorithm are set as follows: the variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\mathbf{e}}$, and $\rho$ obtained by fitting the SEM-Gau model via MLE method. The initial value of the transformation parameter $\gamma$ is set to 1. All fixed-effect parameters of the missing value model ($\boldsymbol{\psi}$) are initialised to 0.1. For the variational covariance matrix, all elements of $\mathbf{B}$ and the diagonal elements of $\mathbf{D}$ are initialised to 0.01. We use $p=4$ factors, as increasing the number of factors did not lead to improved results. Similar initial values are used for the HMC algorithm. Both the HVB-NoB and HMC algorithms are run for 10,000 iterations. Convergence diagnostics plots are provided in Section S9.1 of the online supplement.
@@
Now, we briefly explain how the posterior sample of model parameters, $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and missing values $\boldsymbol{y}_{u}$ are generated in the HVB-NoB method. Once the algorithm converges, the set of variational parameters, $\boldsymbol{\lambda}_{s_{m}}=\left(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top}\right)^{\top}$, is obtained. Given these variational parameters, $10,000$ draws from the variational distribution, $q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\sim N\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\boldsymbol{\mu}_{s_{m}},\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2}\right)$, are generated, which form the posterior sample of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. Based on this posterior sample of model parameters, the posterior sample of $\boldsymbol{y}_{u}^{(i)}$ is generated from the conditional distribution $p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi}^{(i)},\boldsymbol{\psi}^{(i)})$, where $\boldsymbol{\xi}^{(i)}$ and $\boldsymbol{\psi}^{(i)}$ are the $i^{th}$ posterior sample for $i=1,\dots,10,000$.
@@
Figure 2 compares the posterior densities of selected model parameters estimated using HVB-NoB and HMC for fitting the YJ-SEM-Gau model on the simulated dataset with missing data. For most model parameters, the posterior densities obtained from both methods are nearly identical. However, the posterior density of $\gamma$ estimated using the HVB-NoB method is slightly overestimated compared to that of the HMC method.
@@
Figure 2: Comparison of posterior densities of selected parameters in the YJ-SEM-Gau model, obtained using the HMC and HVB-NoB methods on the simulated dataset with missing values. The vertical line represents the true parameter value.
@@
Notably, the computation time per iteration of the VB and HVB-NoB algorithms is significantly lower than that of the HMC algorithm. The computation time per iteration for the VB algorithm is approximately 0.0361, while for the HMC algorithm, it is 0.4680 when estimating the YJ-SEM-Gau model without missing data. Similarly, the computation time per iteration for the HVB-NoB algorithm is approximately 0.0542 seconds, while for the HMC algorithm, it is 41.1658 seconds for estimating the YJ-SEM-Gau model with missing data. The high computational cost of the HMC algorithm in estimating SEMs (such as SEM-Gau) with missing data is also highlighted in the findings of Wijayawardhana et al., (2025).
4.2 Evaluating the performance of proposed SEMs
@@
This section focuses on evaluating the robustness of the proposed SEMs and the VB methods using moderately large simulated datasets. We use a spatial weight matrix from the spData R package (Bivand et al.,, 2023), based on $25,357$ houses sold between 1993 and 1998 in Lucas County, Ohio, USA. Further details on this dataset are provided in Section 5.
@@
We simulate two datasets from the YJ-SEM-t model using the spatial weight matrix $\textbf{W}_{1998}$, which corresponds to 4,378 houses sold in 1998 in Lucas County, Ohio, USA. In the first dataset, we set the parameters $\gamma=0.5$ (right skew) and $\nu=4$ (heavy tails). In the second dataset, we set the parameters $\gamma=1$ (no skew) and $\nu=30$ (light tails). The other parameters are set as follows: 6 fixed effects ($\boldsymbol{\beta}$) are randomly drawn from the discrete uniform values between -3 and 3 (excluding 0), $\sigma^{2}_{\textbf{e}}=0.5$, $\rho=0.8$, and all covariates are generated from a standard normal distribution, $N(0,1)$. Figure 3 shows kernel density plots of the response variables for both datasets.
@@
Figure 3: Kernel density plots of the simulated data: The left panel shows the kernel density of the response variable from YJ-SEM-t with $\nu=4$ and $\gamma=0.5$, displaying right skewness and heavy tails (referred to as dataset 1). The right panel shows the kernel density of the response variable from YJ-SEM-t with $\nu=30$ and $\gamma=1$, which appears nearly symmetric with light tails (referred to as dataset 2).
@@
Section S6 of the online supplement fits the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t models to simulated datasets 1 and 2 without missing data, and evaluates how effectively these models, along with their VB estimation methods, capture the characteristics of these datasets. Section 4.2.1 introduces missing values into both datasets and evaluates the performance of the models and estimation method (the HVB method) in the presence of missing values.
4.2.1 SEMs with missing data

### v2

@@ 5.1 Model comparison with full data
$\text{DIC}_{1}=-4\mathbb{E}_{\boldsymbol{\phi}}[\text{log}~p(\textbf{y}\mid\boldsymbol{\phi})]+2~\text{log}~p(\textbf{y}\mid\overline{\boldsymbol{\phi}}),$ (21)
[…]
@@
[…]
where $\widehat{\boldsymbol{\phi}}$ is the value of $\boldsymbol{\phi}$} that maximises the function $p(\textbf{y}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})$, estimated using samples drawn from the approximate posterior obtained by the proposed VB method. The expectation is also computed using the samples from the approximate posterior.
5.2 Model comparison with missing responses
@@
[…]
where $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})$ is the likelihood of y and m, conditional on model parameters $\boldsymbol{\phi}$ and $\boldsymbol{\psi}$. The values $\widehat{\boldsymbol{\phi}}$, $\widehat{\boldsymbol{\psi}}$, and $\widehat{\textbf{y}}_{u}$ are the posterior samples of $\boldsymbol{\phi}$, $\boldsymbol{\psi}$ and $\textbf{y}_{u}$ that maximise the function $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})p(\boldsymbol{\phi},\boldsymbol{\psi})$, where the samples are drawn from the approximate posterior distribution estimated by the proposed HVB method in Section 4.2. The expectation is also computed using these variational posterior samples. Note that $p(\textbf{y},\textbf{m}\mid\boldsymbol{\phi},\boldsymbol{\psi})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$; see the selection model factorisation in Equation (13) in Section 3.2.
[…]
6 Simulation study: Assessing the accuracy of VB methods
This section compares the accuracy of the proposed VB methods (with full data and with missing data) against the Hamiltonian Monte Carlo (HMC) (Neal,, 2011) method. The HMC method is implemented using the RStan interface (Stan Development Team,, 2023), specifically employing the No-U-Turn Sampler (NUTS) algorithm (Hoffman and Gelman,, 2014). The posterior densities estimated using HMC are considered as ground truth to assess the accuracy of the VB methods. A relatively small simulated dataset ($n=625$) is used for this comparison, as HMC becomes computationally expensive for SAR models with larger datasets, particularly in the presence of missing data (Wijayawardhana et al.,, 2025). The simulation study presented in Section 7 and the real data application in Section 8 use moderately large datasets (n=4,378n=4{,}378).
The prior distributions used for both the VB and HMC algorithms throughout this paper are described next. To map the parameters $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$ onto the real line, we apply the following transformations: $\omega^{\prime}=\text{log}(\sigma^{2}_{\textbf{e}})$, $\rho^{\prime}=\text{log}(1+\rho)-\text{log}(1-\rho)$, $\nu^{\prime}=\text{log}(\nu-3)$, and $\gamma^{\prime}=\text{log}(\gamma)-\text{log}(2-\gamma)$. The prior distributions are then specified on the transformed parameter space. Table 3 summarises the priors and hyperparameters. We use diffuse (non-informative) priors for all model parameters so that posterior inference is driven primarily by the data. For example, the fixed effects $\boldsymbol{\beta}$ are assigned weakly informative normal priors, $\boldsymbol{\beta}\sim N(\textbf{0},100\textbf{I})$, where I is an identity matrix of appropriate dimension (see Table 3). For the relatively large datasets analysed in this paper, the choice of these non-informative hyperparameters has little impact on the parameter estimates or on the convergence behaviour of the VB and HMC algorithms.
Table 3: Prior distributions of model parameters
@@ 6 Simulation study: Assessing the accuracy of VB methods
The comparison in this section is based on a simulated dataset generated from the YJ-SEM-Gau. The dataset contains 625 observations with five covariates, all generated from standard normal distributions. The spatial weight matrix, W, is based on a $25\times 25$ regular lattice, where neighbouring units are defined using the rook neighbourhood (Lloyd,, 2010). The model parameters are specified as follows: the fixed effects ($\boldsymbol{\beta}$) are randomly drawn from a discrete uniform distribution over 3-3 to $3$ (excluding 0). The error variance is set to $\sigma^{2}_{\mathbf{e}}=1$ and the spatial autocorrelation parameter is $\rho=0.8$. The Yeo–Johnson parameter is set to $\gamma=1.25$, introducing moderate left skewness into the simulated response variable. See Figure S1 in Section S7 of the online supplement for the kernel density plot of the response variable. Section 6.1 compares the proposed VB method with HMC using this dataset without missing values, while Section 6.2 compares the proposed HVB method with HMC in the presence of missing responses by introducing missing values into this dataset.
6.1 SEMs with full data
@@
We now assess the accuracy of the VB method for SEMs with full data. The VB algorithm, outlined in Algorithm S1 in Section S2 of the online supplement, and the HMC algorithm are used to fit the YJ-SEM-Gau to the simulated dataset. The initial values for the VB algorithm are set as follows: the variational mean vector ($\boldsymbol{\mu}$) is initialised using the estimates of $\boldsymbol{\beta}$, $\sigma^{2}_{\textbf{e}}$, and $\rho$ obtained from fitting the SEM-Gau via maximum likelihood (ML) estimation method. The variational mean corresponding to the parameter $\gamma$ is initialised to 1. For the variational covariance matrix parameters, the elements of B and the diagonal elements of D are all initialised to 0.01. A value of $p=4$ factors is used. The results do not improve when the number of factors is increased. The HMC algorithm is initialised in the same way, with its initial parameter values set to the initial variational means described above. Both the VB and HMC algorithms are run for 10,000 iterations, at which point convergence is achieved. For the HMC algorithm, the first 5,000 iterations are treated as burn-in and discarded. Convergence for the VB algorithm is assessed visually by inspecting the plots of variational means over iterations. For the HMC algorithms, convergence is assessed by inspecting trace plots of the model parameters (see Section S11.1 of the online supplement).
@@
The HMC algorithm directly generates posterior samples of the model parameters. In contrast, the VB algorithm first estimates the variational parameters, $\boldsymbol{\lambda}$, by running the variational optimisation. Then, 10,000 posterior draws of the model parameters, $\boldsymbol{\xi}=\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},{\sigma}^{2}_{e},\rho,\gamma)^{\top}$}, are generated from the variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi})$; see Section S2 of the online supplement.
Figure 1 compares the posterior densities of selected model parameters estimated using the VB and HMC algorithms. For all model parameters, the posterior densities obtained from both methods are almost indistinguishable and successfully recover the true values.
@@
Figure 1: Comparison of posterior densities of selected parameters in the YJ-SEM-Gau, obtained using the HMC and VB methods on the simulated dataset without missing values. The vertical line represents the true parameter value.
6.2 SEMs with missing responses
@@
Now, missing values are introduced into the response variable y under the MNAR mechanism using the logistic model described in Equation (11), in order to compare the performance of the HVB algorithm presented in Section 4.2 with HMC. To serve as a covariate in the logistic model, an additional variable $\textbf{x}^{\ast}$ is generated from a log-normal distribution with log-scale mean 0 and log-scale standard deviation 1. This simulated covariate, and the response variable y from the YJ-SEM-Gau, are used as the predictors in the missing value model defined in Equation (11). The coefficients are set as follows: $\psi_{0}=-1.0$ (intercept), $\psi_{\textbf{x}^{\ast}}=0.5$ (coefficient of $\textbf{x}^{\ast}$), and $\psi_{\textbf{y}}=-0.1$ (coefficient of y). This configuration yields approximately $50\%$ missing values in y, corresponding to $n_{u}=320$.
[…]
@@
Now, we briefly explain how posterior samples of the model parameters, $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and the missing values $\boldsymbol{y}_{u}$, are generated in the HVB-NoB method. Once the algorithm converges, the set of variational parameters $\boldsymbol{\lambda}=\left(\boldsymbol{\mu}^{\top},\mathrm{vech}(\mathbf{B})^{\top},\mathbf{d}^{\top}\right)^{\top}$ is obtained. Given these variational parameters, 10,000 draws from the variational distribution
$q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})=N\!\left((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top};\,\boldsymbol{\mu},\,\mathbf{B}\mathbf{B}^{\top}+\mathbf{D}^{2}\right)$
are generated, forming the posterior samples of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$. Conditional on these parameter samples, posterior samples of $\mathbf{y}_{u}^{(i)}$ are drawn from
$p(\mathbf{y}_{u}\mid\mathbf{y}_{o},\mathbf{m},\boldsymbol{\xi}^{(i)},\boldsymbol{\psi}^{(i)}),$
using the MCMC steps described in Algorithm 2, where $\boldsymbol{\xi}^{(i)}$ and $\boldsymbol{\psi}^{(i)}$ denote the $i^{\text{th}}$ posterior draws of the model parameters, for $i=1,\dots,10{,}000$.
@@
Figure 2 compares the posterior densities of selected model parameters from the YJ-SEM-Gau and the missing data model, estimated using HVB-NoB and HMC on the simulated dataset with missing values. For most parameters, the posterior densities produced by the two methods are nearly identical. The posterior density of $\gamma$ estimated using HVB-NoB is slightly different compared to HMC, which is expected since HVB is an approximate inference method, whereas HMC yields asymptotically exact posterior samples. Importantly, this minor loss in accuracy is offset by the substantial computational efficiency gained with HVB-NoB, as discussed in Section 6.3.
@@
Figure 2: Comparison of posterior densities for selected parameters of the YJ-SEM-Gau and missing data models, estimated using the HMC and HVB-NoB methods on the simulated dataset with missing values. The vertical line represents the true parameter value.
6.3 Summary and computation time comparison
Overall, the simulation results demonstrate that the proposed VB methods, both with full data and with missing data, provide accurate approximations to the posterior distributions obtained via HMC. For all key model parameters, the VB-based posterior densities closely match those from HMC, with only a slight overestimation observed for the transformation parameter $\gamma$ in the missing data setting.
@@ 6.3 Summary and computation time comparison
Notably, the computation times per iteration of the VB and HVB-NoB algorithms are significantly lower than those of the HMC algorithm. The computation time per iteration for the VB algorithm is approximately 0.0361 seconds, while for the HMC algorithm, it is 0.4680 seconds when estimating the YJ-SEM-Gau without missing data. The computation time per iteration for the HVB-NoB algorithm is approximately 0.0542 seconds, while for the HMC algorithm, it is 41.1658 seconds for estimating the YJ-SEM-Gau with missing data. These results demonstrate that the proposed VB and HVB-NoB methods provide accurate parameter estimates while offering substantial computational savings relative to HMC.
7 Simulation study: Assessing the accuracy and robustness of proposed SEMs
This section evaluates the robustness and accuracy of the proposed SEMs and the VB methods using moderately large simulated datasets. The objective of this simulation study is to assess how well the proposed SEMs capture key characteristics of non-Gaussian spatial data, such as skewness and heavy-tailed distributions, both with full data and in the presence of missing values.
@@ 7 Simulation study: Assessing the accuracy and robustness of proposed SEMs
We generate datasets using a spatial weight matrix from the spData R package (Bivand et al.,, 2023), based on $25,357$ houses sold between 1993 and 1998 in Lucas County, Ohio, USA (see Section 8 for further details). We simulate two datasets from the YJ-SEM-t using the spatial weight matrix $\textbf{W}_{1998}$, corresponding to 4,378 houses sold in 1998:
  1. 1.
    Dataset 1: $\gamma=0.5$ (right skew) and $\nu=4$ (heavy tails).
  2. 2.
    Dataset 2: $\gamma=1$ (no skew) and $\nu=30$ (light tails).
@@
The remaining parameters for both datasets are set as follows: six fixed effect parameters ($\boldsymbol{\beta}$) are randomly drawn from a discrete uniform distribution between -3 and 3 (excluding 0), $\sigma^{2}_{\textbf{e}}=0.5$, $\rho=0.8$, and all covariates are generated from a standard normal distribution, $N(0,1)$. Figure 3 displays kernel density plots of the response variables for both datasets.
@@
Figure 3: Kernel density plots of the simulated data: The left panel displays the kernel density of the response variable from the YJ-SEM-t with $\nu=4$ and $\gamma=0.5$, displaying right skewness and heavy tails (dataset 1). The right panel displays the kernel density of the response variable from the YJ-SEM-t with $\nu=30$ and $\gamma=1$, which appears nearly symmetric with light tails (dataset 2).
@@
Section S8 of the online supplement fits the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t to simulated datasets 1 and 2 with complete data and evaluates how effectively these models, along with their VB estimation methods, capture the underlying characteristics of the datasets. In Section 7.1, missing values are introduced, and the performance of the SEMs along with the HVB estimation method (described in Section 4.2) is assessed under missing data.
7.1 SEMs with missing responses